# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ 1.1.6. Arithmetic progressions in sets given by Piatetski-Shapiro sequences
For $\alpha>1$ and $\alpha\notin\mathbb{N}$, the sequence $(\lfloor n^{\alpha}\rfloor)_{n=1}^{\infty}$ is called Piatetski-Shapiro’s sequence [29], after Piatetski-Shapiro [29] who proved that for all $111<\alpha<12/$ 11 the sequences contain infinitely many prime numbers. In spite of a vast amount of literature, prime numbers in Piatetski-Shapiro’s sequences are not a topic of this paper. There is no integer $\alpha>1$ such that ${\rm PS}(\alpha)$ contains arithmetic progressions of arbitrary lengths [3, 5, 6]. Since ${\rm PS}(\alpha)$ has a convergent reciprocal sum, it is independent of the conjecture of Erdős and Turán.
Piatetski-Shapiro’s sequence is obtained by curving with bounded errors the sequence $(n^{\alpha})_{n=1}^{\infty}$ whose graph $\{(n,n^{\alpha})\in\mathbb{R}^{2}\colon n\in\mathbb{N}\}$ is contained in the graph of the smooth function $x\in[1,\infty)\mapsto x^{\alpha}$. Positive integer sequences with this kind of properties are called slightly curved sequences [32]. Saito and Yoshida [32, Theorem 4] extended Szemerédi’s theorem to graphs of a class of slightly curved sequences, and as a corollary proved that for every $21<\alpha<$ 2, the graph $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in\mathbb{N}\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$ (see [32, Corollary 6]). From this it follows that ${\rm PS}(\alpha)$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$. In fact, the next theorem follows from combining the results in [32]. For completeness we include a proof in Appendix B.
@@
Let $21<\alpha<$ 2. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then $A$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.
Clearly ${\rm PS}(\alpha)$ has polynomial density with exponent $\alpha$ and $\beta=0$. Transferring Theorem 1.10 to the continued fraction via Theorem 1.4 we obtain the following statement.
@@
Let $21<\alpha<$ 2. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+m,k+2m,\ldots,k+\ell m\in A$ and $a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m\ldots,a_{n_{\ell}}(x)=k+\ell m$ for any $x\in E_{{\rm PS}(\alpha),A}.$
@@ 1.1.7. Polynomial progressions in subsets of $\mathbb{N}$ with positive lower density
$\underline{d}(S)=\liminf_{N\to\infty}\frac{1}{N}\#({S\cap[1,N]}).$
@@
Let $S\subset\mathbb{N}$ satisfy $\underline{d}(S)>0$. If $A\subset S$ and $\overline{d}(A|S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{S,A}$.
@@ 1.2. Essential role of non-monotonicity of digits
Arithmetic progressions in the continued fraction was first studied by Tong and Wang [38], who considered the smaller set
$J=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)<a_{n+1}(x)\ \text{ for every }n\geq 1\}\subset E,$
@@
Since the map $x\in J\mapsto\{a_{n}(x)\colon n\in\mathbb{N}\}\in 2^{\mathbb{N}}$ induces a bijection between $J$ and the collection of infinite subsets of $\mathbb{N}$, $J$ is indeed a natural set to investigate. However, the next theorem of independent interest, to be proved in §$\S$4.3, suggests that $E$ rather than $J$ is the right choice for our purpose of building a bridge from density combinatorics to dimension theory of continued fractions.
@@
    If $\overline{d}(S)>0$, then we have
    $\dim_{\rm H}\left\{x\in J\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)\right\}=0<\dim_{\rm H}J.$
@@
    If $S$ has polynomial density with exponent $\alpha<2$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ we have
@@
    dimH{ :xJ:xJ
In sharp contrast to (1.3), Theorem 1.13(a) asserts that for any $S\subset\mathbb{N}$ with positive upper density, the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ forces a significant dimension drop in $J$. In contrast to (1.8), Theorem 1.13(b) asserts that the conditions $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$ and $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)=\overline{d}(A|S)$ together force a dimension drop in $J$. The Hausdorff dimension drops to zero in the case $\alpha=1$, such as $S=\mathbb{P}$ and $S=\mathbb{P}(1)$. These two comparisons show that the non-monotonicity of continued fraction digits for points in $E\setminus J$ is crucial in both Theorems 1.1 and 1.4.
@@ 1.3. Outline of proofs of the main results
Notice that Theorem 1.1(a) follows from Theorem 1.4 since $\overline{d}(S|\mathbb{N})=\overline{d}(S)$ holds for any $S\subset\mathbb{N}$. Moreover, a proof of Theorem 1.1(b) and that of Theorem 1.4 rely on the same set of ideas. Hence, we only outline the proof of Theorem 1.4 below.
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha\geq 1$, and let $A\subset S$ have positive upper density relative to $S$. The construction of the set $E_{S,A}$ in Theorem 1.4 breaks into two main steps. The central idea is to use Moran-like fractals. A Moran fractal is a Cantor-like set with a tractable structure, and has been used extensively for estimating Hausdorff dimension, see [22, 24, 27, 28] for example. In the first step we choose an infinite subset $S_{*}$ of $S$, and construct a Moran-like fractal of Hausdorff dimension $1/(2\alpha)$ consisting of points whose digit sets are contained in $S_{*}$. This set is our seed set, with no relation to $A$. The second step is to modify the seed set into another Moran-like fractal with exactly the same Hausdorff dimension, by adding elements of $A$ into the digit sequences of points in the seed set. The resultant Moran-like fractal is our target set $E_{S,A}$.
@@
The two-step construction of $E_{S,A}$ has been inspired by the afore-mentioned papers [38, 42] that construct subsets of $J$ of Hausdorff dimension $1/2$ corresponding to arithmetic progressions of arbitrary lengths. As already illustrated in §$\S$1.2, an introduction of new ideas is necessary to prove our main results. Since the goal is to construct a subset of $E$, overlapping digits must be avoided. This means that elements of $A$ already appearing in the digit sequences of points in the seed set cannot be added. In order to avoid overlaps of digits, what we actually do in the second step is to add elements of $A\setminus S_{*}$. According to (1.8), the elements of $A$ must appear in the digit set of any point in $E_{S,A}$ with positive relative upper density equal to $\overline{d}(A|S)$. Hence, the upper density of $S_{*}$ relative to $S$ must be zero. This restriction requires an elaboration on the choice of $S_{*}$ in the first step. As such, the seed set we construct in §$\S$3.1 is of maximal Hausdorff dimension and least relative density of digits, called an extreme seed set.
@@
Our main task in the second step is to add sufficiently many elements of $A\setminus S_{*}$ in appropriate positions to achieve the positive density $\overline{d}(A|S)$ without losing Hausdorff dimension of the seed set at all. This seemingly contradictory task can be accomplished by a sharp estimate of Hölder exponent of the associated elimination map, which essentially relies on the non-monotonicity of continued fraction digits for points in $E\setminus J$. See §$\S$3.2 for a general setup of the second step and relevant definitions, and §$\S$3.3 for the key estimate on the elimination map.
@@
The rest of this paper consists of three sections and three appendices. In §$\S 2$ we summarize preliminary results needed for the proofs of the main results and Theorem 1.13. In §$\S 3$ we complete the first step of the construction outlined in §$\S$1.3, and then lay a ground work for the second step. In §$\S 4$ we perform the second step and complete the proofs of the main results. We also complete a proof of Theorem 1.13 in §$\S 4$. In Appendices A and B we prove Theorems 1.2 and 1.10 respectively. In Appendix C we leave comments on generalizations of our main results to a wide class of infinite Iterated Function Systems.
@@ 2. Preliminaries
This section summarizes preliminary results needed for the proofs of the main results and Theorem 1.13. In §$\S$2.1 we recall basic properties of continued fraction digits in the expansion (1.1). In §$\S$2.2 we introduce a Moran fractal, and provide two lower bounds of Hausdorff dimension of sets in $[0,1]$. In §$\S$2.3 we introduce a convergence exponent for infinite subsets of $\mathbb{N}$ and investigate its properties.
2.1. Basic properties of continued fraction digits
For each $n\in\mathbb{N},$ define non-negative integers $p_{n}$ and $q_{n}$ by the recursion formulas
$\begin{split}p_{-1}=1,\ p_{0}=0,\ p_{i}&=a_{i}p_{i-1}+p_{i-2}\ \text{ for }1\leq i\leq n,\\ q_{-1}=0,\ q_{0}=1,\ q_{i}&=a_{i}q_{i-1}+q_{i-2}\ \text{ for }1\leq i\leq n.\end{split}$ (2.1)
For $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, we call the interval
$I(a_{1},\ldots,a_{n})=\begin{cases}\vspace{1mm}\left[\displaystyle{\frac{p_{n}}{q_{n}}},\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}\right)&\text{if}\ n\ \text{is even,}\\ \displaystyle{\left(\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}},\frac{p_{n}}{q_{n}}\right]}&\text{if}\ n\ \text{is odd,}\end{cases}$
an $n$-th fundamental interval. This interval represents the set of elements of $(0,1]$ which have the finite or infinite regular continued fraction expansion beginning by $a_{1},\ldots,a_{n}$, i.e.,
$I(a_{1},\ldots,a_{n})=\{x\in(0,1]\colon a_{i}(x)=a_{i}\ \text{ for }1\leq i\leq n\}.$
The next lemma estimates the length of each fundamental interval in terms of the set of digits prescribing it. Let ||$\cdot$|| ⋅ | denote the Lebesgue measure on $[0,1]$.

### v2

@@ 1.1.6. Arithmetic progressions in sets given by Piatetski-Shapiro sequences
For $\alpha>1$ and $\alpha\notin\mathbb{N}$, the sequence $(\lfloor n^{\alpha}\rfloor)_{n=1}^{\infty}$ is called Piatetski-Shapiro’s sequence [33], after Piatetski-Shapiro [33] who proved that for all $1<\alpha<12/11$ the sequences contain infinitely many prime numbers. Rivat and Wu [36] improved this range to $1<\alpha<243/205$. In spite of a vast amount of literature, the existence of prime numbers in Piatetski-Shapiro’s sequences is not a topic of this paper. There is no integer $\alpha>1$ such that ${\rm PS}(\alpha)$ contains arithmetic progressions of arbitrary lengths [3, 5, 6]. Since ${\rm PS}(\alpha)$ has a convergent reciprocal sum, it is independent of the conjecture of Erdős and Turán.
Figure 1. The graph of a smooth function $x\in[1,\infty)\mapsto f(x)\in\mathbb{R}$ and an associated slightly curved sequence $\{b_{n}\}_{n=1}^{\infty}$: $|b_{n}-f(n)|=O(1)$ $(n\to\infty)$.
Piatetski-Shapiro’s sequence is obtained by curving with bounded errors the sequence $(n^{\alpha})_{n=1}^{\infty}$ whose graph $\{(n,n^{\alpha})\in\mathbb{R}^{2}\colon n\in\mathbb{N}\}$ is contained in the graph of the smooth function $x\in[1,\infty)\mapsto x^{\alpha}$. Positive integer sequences with this kind of properties are called slightly curved sequences [37], see Figure 1. Saito and Yoshida [37, Theorem 4] extended Szemerédi’s theorem to graphs of a class of slightly curved sequences, and as a corollary proved that for every $1<\alpha<2$, the graph $\{(n,\lfloor n^{\alpha}\rfloor)\in\mathbb{N}^{2}\colon n\in\mathbb{N}\}$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$ (see [37, Corollary 6]). From this it follows that ${\rm PS}(\alpha)$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$. In fact, the next theorem follows from combining results in [37]. For completeness we include a proof in Appendix B.
@@
Let $1<\alpha<2$. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then $A$ contains an $\ell$-term arithmetic progression for every $\ell\geq 3$.
Clearly, ${\rm PS}(\alpha)$ has polynomial density with exponent $\alpha$ and $\beta=0$. Transferring Theorem 1.10 to the continued fraction via Theorem 1.4 we obtain the following statement.
@@
Let $1<\alpha<2$. If $A\subset{\rm PS}(\alpha)$ and $\overline{d}(A|{\rm PS}(\alpha))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that
$k+m,k+2m,\ldots,k+\ell m\in A,$
and for any $x\in E_{{\rm PS}(\alpha),A}$,
$a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m\ldots,a_{n_{\ell}}(x)=k+\ell m.$
@@ 1.1.7. Polynomial progressions in subsets of $\mathbb{N}$ with positive lower density
$\underline{d}(S)=\liminf_{N\to\infty}\frac{\#({S\cap[1,N]})}{N}.$
@@
Let $S\subset\mathbb{N}$ satisfy $\underline{d}(S)>0$. If $A\subset S$ and $\overline{d}(A|S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that
$k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S,$
and for any $x\in E_{S,A}$,
$a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m).$
@@ 1.2. Essential role of non-monotonicity of digits
Arithmetic progressions in the continued fraction was first studied by Tong and Wang [43], who considered the set
$J=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)<a_{n+1}(x)\ \text{ for every }n\geq 1\}\subsetneq E,$
@@
Since the map $x\in J\mapsto\{a_{n}(x)\colon n\in\mathbb{N}\}\in 2^{\mathbb{N}}$ induces a bijection between $J$ and the collection of infinite subsets of $\mathbb{N}$, $J$ is indeed a natural set to investigate. However, the next dimension drop theorem of independent interest suggests that $E$ rather than $J$ is a right choice for our purpose of building a bridge from density combinatorics to fractal dimension theory of continued fractions.
@@
    We have
    $\dim_{\rm H}\left\{x\in J\colon\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)>0\right\}=0<\dim_{\rm H}J.$
@@
    If $S$ has polynomial density with exponent $\alpha<2$, then for any $A\subset S$ we have
@@
    dimH{ :xJ
In a sharp contrast to (1.3), Theorem 1.13(a) asserts that for any $S\subset\mathbb{N}$ with positive upper density, the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ forces a significant dimension drop in $J$. In contrast to (1.8), Theorem 1.13(b) asserts that the conditions $\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S$ and $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)=\overline{d}(A|S)$ together force a dimension drop in $J$. The Hausdorff dimension drops to zero in the case $\alpha=1$ such as $S=\mathbb{P}$ and $S=\mathbb{P}(1)$. These comparisons show that the non-monotonicity of continued fraction digits for points in $E\setminus J$ is crucial in both Theorems 1.1 and 1.4. For more technical details on the role of non-monotonicity of digits, see the comments in (III) in Section 1.3.
@@ 1.3. Outline of proofs of the main results
Since $\overline{d}(S|\mathbb{N})=\overline{d}(S)$ holds for any $S\subset\mathbb{N}$, Theorem 1.1(a) follows from Theorem 1.4. Proofs of Theorems 1.1(b) and 1.4 rely on essentially the same set of ideas, but there are some differences. Below we outline proofs of these two.
In principle, constructions of all our target sets in $E$ break into three steps.
  • Step 1: (Section 3.2) construction of an infinite subset of $\mathbb{N}$.
  • Step 2: (Section 3.3) construction of a seed set: a certain Moran fractal in $E$ consisting of points whose digit sets are contained in the set constructed in Step 1.
  • Step 3: (Sections 3.4, 3.5) modification of the seed set constructed in Step 2 without changing the Hausdorff dimension.
Step 3 is most technical. A Moran fractal is a Cantor-like set that is constructed by iterative procedure based on countably many parameters, and has been used extensively for estimating Hausdorff dimension of interesting sets, see [26, 28, 31, 32] for example.
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha\geq 1$, and let $A\subset S$ have positive upper density relative to $S$. The three steps in the construction of the target set $E_{S,A}$ in Theorem 1.4 are described as follows.
  • Step 1: construction of an infinite subset $S_{*}$ of $S$ with zero upper density relative to $S$.
  • Step 2: construction of a Moran fractal $R_{t,L}(S_{*})$ (see (3.1)) of Hausdorff dimension $1/(2\alpha)$ consisting of points whose digit sets are contained in $S_{*}$. This set is our seed set.
  • Step 3: modification of the seed set $R_{t,L}(S_{*})$ into a new Moran fractal of the same Hausdorff dimension, by adding elements of $A$ into the digit sequences of points in the seed set. This new Moran fractal is our target set $E_{S,A}$ (see (4.11)).
This ‘seed set $\rightsquigarrow$ target set construction’ has been inspired by the afore-mentioned papers [43, 47] in which the monotonicity of digits is crucial. However, as pointed out in Section 1.2, for proofs of our main results it is essential to treat points in $E$ for which the digit sequences are not monotone. To this end, we have developed a collection of new ideas summarized as follows.
  • (I)
@@
    Since the goal is to construct a subset of $E$, overlapping digits must be avoided. This means that, in the construction of $E_{S,A}$, elements of $A$ already appearing in the digit sequences of points in the seed set cannot be added in Step 3. In other words, what we actually do in Step 3 is to add elements of $A\setminus S_{*}$ rather than $A$. According to (1.8), the elements of $A$ must appear in the digit set of any point in $E_{S,A}$ with relative upper density $\overline{d}(A|S)$. Hence, the upper density of $S_{*}$ relative to $S$ must be zero. Due to this restriction, the construction of $S_{*}$ in Step 1 is elaborated.
  • (II)
    By the well-known formula in [8, pp.71–72], the Hausdorff dimension of a Moran fractal in $[0,1]$ is estimated from below in terms of their parameters. The seed set we construct in Step 2 and modify in Step 3 consists of points whose digit sets are contained in the set $S_{*}$ constructed in Step 1. Due to this restriction of digits, detecting parameters of our seed set becomes delicate. We do this parameter detection in a unified way, incorporating the case where $S$ has polynomial density with exponent $\alpha\geq 1$. As a by-product, our construction yields a refinement of Ramharter’s theorem [35] of independent interest (see Proposition 3.4).
  • (III)
@@
    Intuitively, adding more digits in Step 3 may result in a more loss of Hausdorff dimension. The task in Step 3 of the construction of $E_{S,A}$ is to add sufficiently many elements of $A\setminus S_{*}$ in appropriate positions, to achieve the positive density $\overline{d}(A|S)$ without losing Hausdorff dimension at all. We accomplish this counter-intuitive task by providing a sharp estimate of Hölder exponent of the associated elimination map. This estimate follows from our careful choices of positions of digits to be added, which relies on the non-monotonicity of continued fraction digits for points in $E\setminus J$. See Section 3.4 for a general setup of Step 3 and relevant definitions, and Section 3.5 for the key estimate on the elimination map.
By virtue of Theorem 1.1(a), in order to prove Theorem 1.1(b) we may start with a set $S\subset\mathbb{N}$ with $\overline{d}(S)=0$ and $d_{\rm B}(S)>0$ (see Section 4.2). The three steps in the construction of the target set $F_{S}$ in Theorem 1.1(b) are described as follows.
  • Step 1: construction of an infinite subset $(\mathbb{N}\setminus S)_{*}$ of $\mathbb{N}\setminus S$ with zero upper density relative to $\mathbb{N}\setminus S$.
  • Step 2: construction of a Moran fractal $R_{t,L}((\mathbb{N}\setminus S)_{*})$ of Hausdorff dimension $1/2$ consisting of points whose digit sets are contained in $(\mathbb{N}\setminus S)_{*}$. This set is our seed set.
  • Step 3: modification of the seed set $R_{t,L}((\mathbb{N}\setminus S)_{*})$ into a new Moran fractal of the same Hausdorff dimension, by adding elements of $S$ into the digit sequences of points in the seed set. This new Moran fractal is our target set $F_{S}$ (see (4.18)).
Due to the difference between the upper density and the upper Banach density, Step 3 of the construction of $F_{S}$ is considerably different from Step 3 of the construction of $E_{S,A}$. We have designed Step 3 in Sections 3.4 and 3.5 with a sufficient generality, to accommodate both constructions in a unified manner.
1.4. Structure of the paper
@@
The rest of this paper consists of four sections and two appendices. In Section 2 we summarize preliminary results. In Section 3 we carry out the construction of a Moran fractal in the three steps outlined in Section 1.3. In Section 4 we apply the general construction in Section 3, and complete the proofs of Theorems 1.1 and 1.4. We also complete a proof of Theorem 1.13. In Section 5, we generalize our main results to a wide class of Iterated Function Systems and semi-regular continued fractions. In Appendices A and B we prove Theorems 1.2 and 1.10 respectively.
@@ 2. Preliminaries
This section summarizes preliminary results needed for the proofs of Theorems 1.1, 1.4 and 1.13. In Section 2.1 we recall basic properties of continued fraction digits in the expansion (1.1). In Section 2.2 we introduce a Moran fractal, and provide two lower bounds of Hausdorff dimension of sets in $[0,1]$. In Section 2.3 we introduce a convergence exponent for infinite subsets of $\mathbb{N}$, and investigate its properties.
2.1. Fundamental intervals for continued fractions
For each $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, define non-negative integers $p_{n}$ and $q_{n}$ by the recursion formulas
$\begin{split}p_{-1}=1,\ p_{0}=0,\ p_{i}&=a_{i}p_{i-1}+p_{i-2}\ \text{ for }i=1,\ldots,n,\\ q_{-1}=0,\ q_{0}=1,\ q_{i}&=a_{i}q_{i-1}+q_{i-2}\ \text{ for }i=1,\ldots,n.\end{split}$ (2.1)
For $n\in\mathbb{N}$ and $(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, we define an $n$-th fundamental interval by
$I(a_{1},\ldots,a_{n})=\begin{cases}\vskip 2.84526pt\left[\displaystyle{\frac{p_{n}}{q_{n}}},\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}}\right)&\text{if}\ n\ \text{is even,}\\ \displaystyle{\left(\frac{p_{n}+p_{n-1}}{q_{n}+q_{n-1}},\frac{p_{n}}{q_{n}}\right]}&\text{if}\ n\ \text{is odd.}\end{cases}$
This interval represents the set of elements of $(0,1]$ which have the finite or infinite regular continued fraction expansion beginning by $a_{1},\ldots,a_{n}$, i.e.,
$I(a_{1},\ldots,a_{n})=\{x\in(0,1]\colon a_{i}(x)=a_{i}\ \text{ for }i=1,\ldots,n\}.$
Let $|\cdot|$ denote the Euclidean diameter of sets in $[0,1]$. The next lemma estimates the diameter of each fundamental interval in terms of its string of digits.